OpenStudy (anonymous):

The center of a circle is at the origin. An endpoint of a diameter of the circle is at (-3, -4). How long is the diameter of the circle? 25?

OpenStudy (lgbasallote):

distance from endpoint ofdiameter to center is \[\sqrt{(-3 - 0)^2+(-4-0)^2} \implies\sqrt{9 + 16} \implies \sqrt{25} \implies 5\] so the radius is equal to 5..what do you think is the diameter then?

OpenStudy (anonymous):

10

OpenStudy (ganpat):

sorry, that was wrong.. :-p

OpenStudy (lgbasallote):

right